OpenStudy (anonymous):

What is the general form of an equation that represents the line that passes through the points ( -9, -5) and (3, 8) ? A) 12x-13y-57=0 B) 13x+12y-57=0 C) 13x-12y+57=0 D) 12x+13y+57=0

OpenStudy (anonymous):

first find the slope of the line. \[slope = m=\frac{y2-y1}{x2-x1}=\frac{8-(-5)}{3-(-9)}=\frac{13}{12}\] \[y = mx+b\] plug in one of the order pairs into the above equation to find b. then just move everything to the left side of the = sign.

OpenStudy (anonymous):

Eh I dont get it..

OpenStudy (anonymous):

It only shows me how to find the slope not the whole equation.

OpenStudy (raden):

use the formula : (y - y1)/(y2-y1) = (x - x1)/(x2 - x1)

OpenStudy (abb0t):

\[(y-y_1)=m(x-x_1)\] then use algebra to rearrange it to standard form: y = mx+b